Shift summary for Owl, 8/2/99

Ed Blucher (blucher@hep.uchicago.edu)
Mon, 2 Aug 1999 06:50:40 -0500

Shift Summary for Owl Shift, 08/2/99

On Shift: EB,VJ

Hours of beam available at all intensities: 6
Hours of beam available at nominal intensity: 6 (~3.9e12)
Hours of E832 data (special runs not included): 5.8
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: run start, KFRENDs (mostly for crate 53)

========================================================================
Runs ended this shift:
Run 14084 - BEAM_832 - 12461 KL -> 2pi0

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
E832 Datataking

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14084 2:55 TeV Refrigerator problems - no estimate.
3:02 Beam returns
4:00 Linac power supply problems - no estimate.
4:50 Stop run and update adc peds.
4:54 Beam returns
14085 4:56 Start run

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

2E12 run will begin during the next shift.